Solve the following linear system.
13y = 20x + 3y
13y + ( 8x + y ) = 180

Answer by FrankM(1040)

13y = 20x + 3y simplifies to
13y-3y=20x
10y=20x
y=2x
second eq
13y + ( 8x + y ) = 180
14y+8x=180
substitute
14(2x)+8x=180
28x+8x=180
36x=180
x=5
y=10